MrQuestion:

A student attempted to generate an equivalent expression using the distributive property, as shown below: 6(m + 3) = 6m + 3 What was the mistake made? 6 was not multiplied by m Incorrect sign was used for the first term Incorrect sign was used for the second term 6 was not multiplied by 3

snowflake0531:

do you know how the distributive property works?

MrQuestion:

No not realy why

snowflake0531:

since this is distributive property a(b+c) = ab+ac a(b-c) = ab-ac

XioGonz:

The answer should be 6(m+3)= 6m+ 18......So what does this tell you

snowflake0531:

do you know what they did wrong now?

MrQuestion:

No thats why i asked

snowflake0531:

but i just told you that a(b+c) = ab +ac also, look at whatxiogonz said

MrQuestion:

I think answer is option 2

snowflake0531:

alas, no

snowflake0531:

a(b+c)=ab +ac so, they did 6(m+3) only ab, not ac, it's supposed to be 18, not 3, so they didn't mutliply the 3 with 6